Two spherical objects have masses of 8000kg and 1500 kg. Their centers are separated by a distance of 1.5m. Find the gravitational attraction between them.


1
Expert's answer
2022-04-28T16:42:22-0400
"F=G\frac{m_1m_2}{r^2}"

"F=\frac{6*67*10^{-11}*8000*1500}{1.5^2}=3.6*10^{-4}\:\rm N"